Lacrue Inc. has provided the following data concerning one of the products in its standard cost system. Variable manufacturing overhead is applied to products on the basis of direct labor-hours.

Inputs Standard Quantity or Hours per Unit of Output Standard Price or Rate
Direct materials 6.3 ounces $ 5.50 per ounce
Direct labor 0.40 hours $ 22.00 per hour
Variable manufacturing overhead 0.40 hours $ 3.80 per hour

The actual output for the period was 3,700 units.

The standard amount of materials allowed for the actual output is closest to:

Garrison 16e Rechecks 2017-07-14

Multiple Choice

  • 22,687 ounces

  • 23,300 ounces

  • 23,310 ounces

  • 23,940 ounces

Abburi Company's manufacturing overhead is 35% of its total conversion costs. If direct labor is $83,200 and if direct materials are $29,700, the manufacturing overhead is:

Garrison 16e Rechecks 2017-08-01

Multiple Choice

  • $154,514

  • $15,992

  • $44,800

  • $60,792

Winder Corporation is a specialty component manufacturer with idle capacity. Management would like to use its extra capacity to generate additional profits. A potential customer has offered to buy 3,000 units of component QEA. Each unit of QEA requires 5 units of material F85 and 5 units of material E71. Data concerning these two materials follow:

Material Units in Stock Original Cost Per Unit Current Market Price Per Unit Disposal Value Per Unit
F85 740 $ 4.90 $ 4.75 $ 4.20
E71 13,680 $ 5.00 $ 4.70 $ 3.60

Material F85 is in use in many of the company's products and is routinely replenished. Material E71 is no longer used by the company in any of its normal products and existing stocks would not be replenished once they are used up.

What would be the relevant cost of the materials, in total, for purposes of determining a minimum acceptable price for the order for product QEA?

Multiple Choice

  • $126,702

  • $126,295

  • $145,965

  • $141,750

Boney Corporation processes sugar beets that it purchases from farmers. Sugar beets are processed in batches. A batch of sugar beets costs $64 to buy from farmers and $23 to crush in the company's plant. Two intermediate products, beet fiber and beet juice, emerge from the crushing process. The beet fiber can be sold as is for $34 or processed further for $37 to make the end product industrial fiber that is sold for $85. The beet juice can be sold as is for $55 or processed further for $41 to make the end product refined sugar that is sold for $85.

What is the financial advantage (disadvantage) for the company from processing one batch of sugar beets into the end products industrial fiber and refined sugar rather than not processing that batch at all?

rev: 12_15_2017_QC_CS-112029

Garrison 16e Rechecks 2017-12-15

Multiple Choice

  • ($4) per batch

  • $21 per batch

  • $5 per batch

  • ($165) per batch

Puello Corporation has provided the following data concerning an investment project that it is considering:

Initial investment $ 480,000
Annual cash flow $ 145,000 per year

Click here to view Exhibit 13B-1 and Exhibit 13B-2, to determine the appropriate discount factor(s) using the tables provided.

The life of the project is 4 years. The company’s discount rate is 8%. The net present value of the project is closest to:

Multiple Choice

  • $240

  • $480,240

  • $100,000

  • $480,000

Which of the following is correct? The break-even point occurs on the CVP graph where:

Multiple Choice

  • total variable expenses equal total contribution margin.

  • total profit equals total fixed expenses.

  • total profit equals total expenses.

  • total contribution margin equals total fixed expenses.
Answer #1

Standard material for actual output = standard material per unit*actual units

6.3*3700

=23,310 ounces

Conversion cost = direct labour plus manufacturing overhead

Conversion cost =83,200/65%

=$128,000

Manufacturing overhead = 128,000*35% =$44,800

Relevant cost of F85 = current market value = 5*4.75*3,000 =71,250

Of E71 = disposal value of those in stock + market value of remaining

= 13,680*3.6 + 1,320*4.70 =$55,452

Total relevant cost

=$126,702

Financial advantage of processing = selling price after processing - total cost

= 85+85-37-41-64-23

=$5 per batch

NPV = present value of cash inflows - present value of cash outflows

= 145,000*PVAF(8%, 4 years) - 480,000

=$240

Total contribution margin equals total fixed expenses